• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuchina

    Log entry time 16:00:34 on April24,2007

    Entry number 199883

    keyword=Day shift summary

    It was a pretty smooth shift.

    There were 2 runs on Dummy with high current of 90uA while the limit was clearly posted to be 40uA. Hopefully, we didn't ruin the dummy target.

    While we were moving the target, Q1 magnet tripped to zero. We had to reset it mannually by pressing the button in the middle room. It took about half an hour to recover from this trip.

    After that we returned to 4He target, beam was not steady for about an hour, and there was no beam for another half an hour (it was taken off for tuning). Otherwise everything was smooth.

    The summary of runs:
    Run Target Events Current
    2732-4 4He 1.25M 90uA
    2735,6 Dummy 90uA High Current!!!
    2738,9 Dummy 40uA
    2740-6 4He 1.25M 90uA